2x-3y=-24
x+6y=18



Answer by oberobic(2304)

2x -3y = -24
x + 6y = 18
.
multiply the first equation by 2 and then add
.
4x -6y = -48
x + 6y = 18
--------------
5x = -30
.
x = -6
.
substitute to find 'y'
.
x + 6y = 18
-6 + 6y = 18
6y = 24
y = 4
.
Use the other equation to check the answer.
2x -3y = 2(-6) -3(4) = -12-12 = -24
Correct.
.
Answer: x = -6 and y = 4
.
Done.